50 tons of Fe2O3 is reacting with 30 tons of C.




find out mass of Fe extracted?




Volume of Co2 gas released in dm3?

1
Expert's answer
2021-11-05T01:07:37-0400

Solution:

The molar mass of Fe2O3 is 159.69 g/mol

The molar mass of C is 12.011 g/mol


Determine moles of each reactant:

(50 tons Fe2O3) × (1000000 g / 1 ton) × (1 mol Fe2O3 / 159.69 g Fe2O3) = 313106.64 mol Fe2O3

(30 tons C) × (1000000 g / 1 ton) × (1 mol C / 12.011 g C) = 2497710.432 mol C


Balanced chemical equation:

2Fe2O3 + 3C → 4Fe + 3CO2

According to the chemical equation above:

2 mol of Fe2O3 reacts with mol of C

Thus, 313106.64 mol of Fe2O3 reacts with:

(313106.64 mol Fe2O3) × (3 mol C / 2 mol Fe2O3) = 469659.96 mol C

However, initially there is 2497710.432 mol of C (according to the task).

Therefore, Fe2O3 acts as limiting reagent and C is excess reagent.


According to stoichiometry:

2 mol of Fe2O3 produces 4 mol of Fe

Thus, 313106.64 mol of Fe2O3 produces:

(313106.64 mol Fe2O3) × (4 mol Fe / 2 mol Fe2O3) = 626213.28 mol Fe


The molar mass of Fe is 55.845 g/mol

Therefore,

(626213.28 mol Fe) × (55.845 g Fe / 1 mol Fe) = 34970880.62 g Fe

(34970880.62 g Fe) × (1 ton / 1000000 g) = 34.97 tons Fe = 35 tons Fe


According to stoichiometry:

2 mol of Fe2O3 produces 3 mol of CO2

Thus, 313106.64 mol of Fe2O3 produces:

(313106.64 mol Fe2O3) × (3 mol CO2 / 2 mol Fe2O3) = 469659.96 mol CO2


At STP, one mole of any gas occupies a volume of 22.4 dm3

Thus, 469659.96 mol of CO2 occupies:

(469659.96 mol CO2) × (22.4 dm3 / 1 mol) = 10520383.1 dm3 CO2 = 10.5×106 dm3 CO2


Answers:

35 tons of Fe

10.5×106 dm3 of CO2